A number is selected at random from the set of 1, 2, 3, 4, 5, 6, 7, 8, 9, and 10. Find:
P(even or power of 3); P(power of 2 or even); P odd and prime); P(even or prime); P(prime give odd)

Answer by stanbon(75887)

A number is selected at random from the set of 1,2,3,4,5,6,7,8,9,10.
Find:
P(even or power of 3) = P(even)+P(pwr of 3) = (5 + 3)/10 = 8/10
---------------------------
P(power of 2 or even) = P(pwr of 2) + P(even)-P(pwr 2 OR even)
= (4+5-3)/10 = 6/10
---------------------------
P (odd and prime) = 3/10
-----------------------------------
P(even or prime) = P(even)+P(prime)-P(even AND prime) = (5+3-1)/10 = 7/10
-----------------------------------
P(prime given odd) = P(prime AND odd)/P(odd) = (3/10)/(5/10) = 3/5
===============================
Cheers,
Stan H.
===============================